# Semi-supervised and unsupervised learning for health indicator extraction from guided waves in aerospace composite structures

### v1

@@ 2 Method
This section provides a detailed overview of the methodology adopted in this study, encompassing the experimental setup and the framework for HI generation. Subsection 2.1 introduces the experimental setup which forms the foundation of the analysis. Subsection 2.2 describes the overall proposed two-stage framework for HI generation, followed by Subsection 2.3, which outlines the criteria used to evaluate the generated HIs. Subsection 2.4 delves into the signal processing techniques applied to preprocess the data, while Subsection 2.5 focuses on the feature extraction process. Finally, Subsection 2.6 presents the feature fusion approach, which include Diversity-DeepSAD, DTC-VAE, and ensemble learning models.
@@ 2.1 Experimental setup and data preparation
To evaluate the proposed framework, five single-stiffener composite panels were subjected to run-to-failure compressive fatigue loading under ReMAP project [^4] ReMAP: Real-time Condition-based Maintenance for Adaptive Aircraft Maintenance Planning. https://h2020-remap.eu/. The panels consist of IM7/8552 carbon fiber-reinforced epoxy in unidirectional prepreg, formed with layups of $\pm 45/0/45/90/-45/0]_{S}$} and $\pm 45/0/\pm 45]_{S}$} [15]. The panels were monitored using different SHM techniques [60], from which we use only GWs in order to construct HIs.
@@
The input features data was arranged in 5 folds, in each of which a different specimen was used for testing while the remaining 4 were for training. Training and test data were normalised relative to the training data only using Z-score scaling, in other words, subtracting by the mean and dividing by the standard deviation.
The model output was also normalised using the training data, but with min-max normalisation, to ensure HIs had a value of approximately 0 in the first timestep (healthy state) and approximately 1 at the last (failure), although this has no effect on fitness scores.
2.2 Two-Stage framework for HI generation
This study follows a two stage framework resulting in HI generation through the use of signal processing techniques and AI models. In the two stages, as can be seen in Figure 2, time domain measurements were processed using different SP methods, including fast Fourier transform (FFT), short-time Fourier transform (STFT), empirical mode decomposition (EMD), and Hilbert transform (HT). Statistical features were then extracted from the processed signals for each path, actuator frequency, and measurement time step. As all input features should be ideally of the same SP method to ensure efficient processing, one SP method can be practically used in the workflow. With this in mind, the two methods with the highest performance were selected for further individual use.
In the third stage, as can be seen in Figure 2, the features selected were fed into AI models (fourth phase), i.e., the Diversity-DeepSAD and DTC-VAE models, where hyperparameters are fine-tuned using Bayesian optimisation. This process was repeated for each combination of excitation frequency of PZTs and SP method. The extracted HIs from each actuating frequency were then fused by an unsupervised ensemble learning model, enabling the evaluation of the remaining combinations of AI model and SP method against the HIs evaluation criteria to determine the optimal final combination.
Figure 2: Framework for HI generation and evaluation — Step 1: Data Pre-Processing, Step 2: Signal Processing, Step 3: Feature Representation, Step 4: Feature Fusion — with prognostic criteria (Mo, Pr, Tr).
@@ 2.3 HIs criteria
Given the lack of true HI labels, evaluating the quality of the developed HIs necessitates leveraging known physics-based principles and facts. The HIs criteria include Mo, Pr, and Tr, and they are standard for this application [15, 23, 61, 13, 46]. These criteria range from 0 to 1, with 1 being the ideal value.
Mo measures the extent to which the HI shows only an increasing or decreasing trend. If maintenance is not performed, health should always decrease with an increasing number of cycles, and thus a monotonic HI should more closely model reality. The modified Mann-Kendall (MMK) formulation of Mo is presented in Equation 1 [15]:
@@
where $M$ is the number of specimens, $N$ is the number of GW measurements (time steps), $y^{m}_{j}$ denotes a HI at timestep $j$ of specimen $m$, $t_{j}$ is the time at the $j^{th}$ measurement, $sgn(\cdot)$ is the sign function. This formulation of Mo is preferable as it considers time gaps greater than one unit, thus reducing the effect of noise and favoring monotonicity through the whole trend, as opposed to between adjacent points [15].
Pr measures the HIs deviation at failure between specimens. Preferably, all specimens should have the same value of HI at the end of life, and HIs can be reliably used to predict failure. Equation 2 expresses Pr [15], where $std(\cdot)$ refers to the standard deviation between specimens.
@@
Finally, Tr is a measure of the correlation between HIs of different units [15]. The set of specimens should show the same HI pattern from beginning to end of life, as this makes their behaviour more predictable. Tr is given by Equation 3, where $corr(\cdot)$ is the Pearson correlation function.
@@
To effectively evaluate a model’s performance on test specimens while mitigating bias towards highly matched training specimens and avoiding the obscuring of unmatched test specimen behavior, alternative evaluation criteria must be established. These criteria should prioritise the behavior of test specimens over those used during training, ensuring a more reliable assessment [15, 24]. These criteria include $Mo_{test}$, which evaluates the monotonicity of only the test specimen(s), and $Pr_{test}$, which assesses the prognosability by emphasizing the deviation of the test specimen(s) relative to those used for training. These metrics are defined in Equation 4 and Equation 5 [15]:
@@
where a superscript traintrain or testtest represents train or test specimens, respectively, and thus $M^{train}$ refers to the number of the training specimens.
Finally, the fitness score $F_{all}$ is defined by Equation 6, and measures the overall quality of HIs [15], while $F_{test}$ is defined similarly in Equation 7 with a focus on test units. Setting the control constants $k_{Mo}$, $k_{Pr}$, and $k_{Tr}$ equal to 1, in order to weight the criteria equally, the maximum fitness score is 3; the fitness may also be given as a percentage relative to this maximum.
@@
Both fitness scores $F_{all}$ and $F_{test}$ are calculated and discussed in this study.
@@ 2.4 Signal processing
Signal processing (SP) techniques were applied to transform the measured signals into the frequency and time-frequency domains, in order to highlight relevant information, such as frequency components or noise isolation. These transformations enabled the extraction of relevant features for further analysis. The techniques used include FFT, STFT, EMD, and HT.
The FFT algorithm transforms the original temporal signal into the frequency domain by processing the entire time domain at once and outputting corresponding amplitudes for the underlying frequencies. However, it does not retain temporal variations in the signal, such as how these frequencies evolve over time, which may reveal additional information about the structure’s health. In contrast, STFT preserves time-frequency information by dividing the data into smaller time intervals and analyzing localised segments, making it advantageous for non-stationary signals where frequency content changes over time. The result of STFT is a 2D matrix, where time and frequency are the axes, and amplitude is the value at each position. Nevertheless, there is a nonlinear trade-off between time and frequency resolution: too many time segments reduce the available frequency samples, and vice versa. In this study, it was found that performing STFT with an FFT length of 250 maximised the number of data points, resulting in an overlap length of 125.
In the case of EMD, the sifting algorithm was implemented to iteratively decompose the signal into intrinsic mode functions (IMFs) in the time domain. The process involves constructing upper and lower envelopes of the signal using cubic splines to interpolate local maxima and minima, respectively. The mean of these envelopes is then subtracted from the signal at each iteration, refining the signal until it satisfies the conditions for an IMF. Suitable stopping conditions are necessary to avoid valuable information being lost [62]. Through trial and error, these criteria were set to a minimum standard deviation of $0.1$ and a maximum of 10 iterations.
The HT algorithm produces a complex-valued signal, the analytic signal, from a real signal. The real component represents the original signal, while the imaginary component is created by phase-shifting each frequency component by $\frac{\pi}{2}$ [63]. The analytic signal therefore contains information concerning instantaneous amplitude and phase.
@@ 2.5 Feature extraction
Statistical features, as detailed in A, were extracted from the time, frequency, and time-frequency domains to characterise the signals comprehensively. Common statistical features for the time and frequency domains were implemented following [23].
@@
For the first stage, features across different actuator-sensor paths and frequencies are averaged, resulting in 139 features per time step (GW measurement) for each specimen. After obtaining all time steps and specimens, the HI prognostic criteria defined in Subsection 2.3 are applied to evaluate each feature. To avoid anomalously poor features disproportionately impacting the SP method selected, the mean fitness score across all features is calculated to serve as a benchmark. Features with fitness scores exceeding this benchmark are retained for further analysis. The SP method yielding features with the highest remaining mean fitness scores is identified as optimal for developing an efficient framework to extract HIs.
@@ 2.6 Feature fusion
Given the extracted and selected features as inputs, two deep learning models were developed to generate HIs, i.e., semi-supervised Diversity-DeepSAD and unsupervised DTC-VAE, which will be presented in this section.
@@ 2.6.1 Diversity-DeepSAD
The original DeepSAD model incorporates labeled training data in the loss function of the otherwise unsupervised deep support vector data description (Deep SVDD) model [42]. This approach transforms input data into a hyperspace of reduced dimensions, in which the distance to a point, the hypersphere centre $\mathbf{c}$, represents the extent to which the datapoint is an anomaly. Frusque et al. [46] modified the loss function further to include a term to increase diversity, resulting in Diversity-DeepSAD for HI generation. This was demonstrated on a rotational machinery dataset with hundreds of measurement timesteps, a condition not often met in many other applications. For example, the ReMAP dataset may contain as few as 30 timesteps before failure. Consequently, it is necessary to explore the impact of alternative functions for generating auxiliary artificial labels and to evaluate the influence of the number of labeled measurements.
In prior studies [42, 46], a binary set of auxiliary labels $\tilde{y}_{j}\in\{1,-1\}$ was employed, where a label of 1 was assigned to measurements representing a very healthy condition, and -1 corresponded to those reflecting a severely damaged state. This approach focuses solely on two extreme conditions. However, these two phases typically occupy only a small portion of the system’s entire lifetime, limiting the potential to fully leverage auxiliary labels. This issue is particularly pronounced in the severely damaged phase, where measurements are rare, often with only a single observation available at the point of failure for assigning the label -1. To address these limitations, this study generated continuous auxiliary labels $\tilde{y}_{j}\subset[1,-1]$ according to Equation 8:
@@
where, for simplicity, a linear function was applied for the first and last quarters of the lifetime. The quarter ratio was established experimentally to produce improved results. As visualised in 3(a), the remaining data points do not leverage auxiliary labels, providing flexibility to the model, given the uncertainty of the true health status for these intermediate conditions.
It should be noted that auxiliary labels $\tilde{y}_{j}$ differ from the expected health indicator (HI), denoted by $y_{j}$. Assuming the hypersphere center $\mathbf{c}$ represents the healthy state, the HI is defined as the distance from that center:
@@
where $z_{j}$ denotes the latent space representation obtained from the autoencoder. When $z_{j}$ is close to the healthy state $\mathbf{c}$, the $y_{j}$ approaches 0, indicating minimal deviation from the healthy condition. Conversely, as $z_{j}$ moves farther from $\mathbf{c}$, the $y_{j}$ increases, reflecting a more degraded condition. In this framework, the maximum value of 1 is assigned to $y_{j}$ in the fully damaged state. Consequently, the HI can freely exhibit linear or non-linear behavior—with different main trends, such as exponential or polynomial behavior—between these bounds, as illustrated in 3(b).
(a) Auxiliary labels generated for training as a function of lifetime. Data collected in the dashed region is left unlabelled.
(b) Examples of possible HI embeddings, where each curve represents a basic function capturing the main trend, either individually or in combination. Note that 0 corresponds to healthy and 1 to failure, unlike auxiliary labels (1 = healthy, -1 = failure).
@@
Denoting the encoder and decoder functions by $\phi$ and $\psi$, respectively, the HI can be expressed as:
@@
where $\nu$ and $\lambda$ are hyperparameters controlling the weights of the regularization terms $\mathcal{L}_{L2}$ and $\mathcal{L}_{D}$, respectively. The regularization terms are defined as follows:
@@
$\mathcal{L}_{D}=\sum_{h=1}^{H}(\zeta_{h}-\ln(\zeta_{h}))$ (14)
where $L$ denotes the number of layers in the model, and $\zeta_{h}$ represents the $h^{\text{th}}$ eigenvalue of the Gram matrix $G=Z^{T}Z$, in which $Z$ is the DeepSAD embedding. $\mathcal{L}_{L2}$ represents $L_{2}$ regularization, minimizing the sum of squares of node weights to encourage a greater distribution of smaller weights and prevent overfitting, while $\mathcal{L}_{D}$ promotes a greater distribution of training data in the model output [46]. These three terms therefore work together to ensure that the model effectively learns the latent representations, while the regularization terms help prevent overfitting and promote diversity within the learned embeddings.
A standard autoencoder (AE) is always used to pretrain the network weights (learnable parameters $\theta$). This AE is formed by adding a decoder $\psi$ which inverts the Diversity-DeepSAD network layers, considering the mean squared difference between input and output data as a loss function, $\mathcal{L}_{AE}$. This loss function is used to pre-train the autoencoder, but is disregarded in the training of Diversity-DeepSAD in Equation 12. The hypersphere center $\mathbf{c}$ is initialised by computing the average output from the model during a forward pass on the training data. Initially, $\mathbf{c}$ is set to a zero matrix (e.g., dimensions $4\times 4$). In the pretraining phase, denoted by $\phi_{\theta}^{[0]}$, the model performs a forward pass on the training data, and the outputs for each sample are accumulated. The number of samples is incremented with each data point, and the cumulative outputs are added to $\mathbf{c}$. After processing all the data, the matrix $\mathbf{c}$ is normalised by the total number of samples to obtain the average. To avoid numerical issues, any component of $\mathbf{c}$ with an absolute value smaller than $\epsilon$ is adjusted as follows:
[…]
This adjustment ensures stability in the hypersphere center values, helping the model establish a reference point for the healthy state. The hypersphere center $\mathbf{c}$ serves as a critical foundation for anomaly detection and health indicator estimation. The model architecture is illustrated in Figure 4, where red, yellow and green represent pretraining, training and model outputs respectively.
@@
Diversity-DeepSAD architecture and hyperparameter optimisation: The Diversity-DeepSAD model was implemented with 6 hidden layers, each containing half the number of neurons as the previous layer, and leaky rectified linear units (Leaky ReLU) used as activation functions between layers. The hyperspace was set to be 16-dimensional (resulting in $\mathbf{c}$ with dimensions $4\times 4$), which was found to enhance the performance of AE during pretraining. Key hyperparameters, including the number of of epochs, learning rate, and batch size—those observed to have a strong impact on the results—were optimised using Bayesian optimisation with Gaussian processes, based on the search space shown in Table 2. Maximizing the $F_{all}$ score has been set as the objective function of the Bayesian optimization algorithm. To mitigate the model’s tendency to separate the failure state too distinctly from intermediate health degradation, a low maximum learning rate was set. All remaining hyperparameters were fixed, with their values determined from literature [46], namely $\nu=10$, $\eta=10$, $\lambda=0.001$ and $\epsilon=1\cdot 10^{-6}$.

### v2

@@ 3 Method
This section provides a detailed overview of the methodology adopted in this study, encompassing the experimental setup and the framework for HI generation. Subsection 2.1 introduces the experimental setup which forms the foundation of the analysis. Subsection 2.2 describes the overall proposed two-stage framework for HI generation, followed by Subsection 2.3, which outlines the criteria used to evaluate the generated HIs. Subsection 2.4 delves into the SP techniques applied to preprocess the data, while Subsection 2.5 focuses on the feature extraction process. Finally, Subsection 2.6 presents the feature fusion approach, including Diversity-DeepSAD, DTC-VAE, and ensemble learning models. Details on code availability are provided in C.
@@ 3.1 Experimental setup and data preparation
To evaluate the proposed framework, five single-stiffener composite specimens were subjected to run-to-failure compressive fatigue loading under the ReMAP-H2020 project [Dataverse]. The specimens consist of IM7/8552 carbon fibre-reinforced epoxy in unidirectional prepreg, formed with layups of $\pm 45/0/45/90/-45/0]_{S}$} and $\pm 45/0/\pm 45]_{S}$} [Moradi2024]. The specimens were monitored using different SHM techniques, from which only GWs are used to construct HIs.
@@
Figure 2 presents example raw GW signals from the initial (left) and end-of-life (right) states of one composite skin–stiffener panel (composite specimen 5, 125 kHz excitation). Each GW measurement consists of 56 actuator–sensor paths recorded at 6 excitation frequencies, resulting in a 56 × 6 × 2000 data tensor per sample. Zoom-in views highlight changes in waveform characteristics across fatigue progression. The bottom plot shows the corresponding HIs predicted across the lifetimes of multiple units, demonstrating both unit-to-unit variability and complete run-to-failure coverage. In this example, unit 5 was held out entirely for testing, while the remaining units were used for training and validation.
Figure 2: Example initial and end-of-life guided wave signals and corresponding HIs illustrating the structure of the dataset and fatigue-induced waveform evolution.
The input feature data were arranged in 5 folds, in each of which a different specimen was used for testing while the remaining 4 were for training. This leave-one-unit-out strategy ensures that all measurements from the held-out panel, including its manufacturing tolerances and experimental uncertainties, remain completely unseen during training, providing a stringent cross-unit generalisation test and helping to mitigate overfitting despite the limited number of structural units. Training and test data were normalised relative to the training data only using Z-score scaling, in other words, subtracting by the mean and dividing by the standard deviation.
The model output was also normalised using the training data, with min-max normalisation, to ensure HIs had a value of approximately 0 in the first timestep (healthy state) and approximately 1 at the last (failure).
3.2 Four-stage framework for HI generation
This study followed a four-stage framework, as can be seen in Figure 3, resulting in HI generation through the use of SP techniques and AI models. The first stage was pre-processing data to the correct format. In the second stage, time domain measurements were processed using different SP methods, including fast Fourier transform (FFT), short-time Fourier transform (STFT), empirical mode decomposition (EMD), and Hilbert transform (HT). Thirdly, statistical features were extracted from the processed signals for each path, actuator frequency $f$, and measurement timestep. As all input features should ideally be of the same SP method to ensure efficient processing, a single SP method is used in the workflow. With this in mind, the two SP methods with the highest performing features were selected for further individual use.
In the fourth stage, the features selected were fed into AI models, i.e. the Diversity-DeepSAD and DTC-VAE models, where hyperparameters were fine-tuned using Bayesian optimisation. This process was repeated for each combination of excitation frequency of PZTs and SP method. The extracted HIs (i.e. $y$) from each actuating frequency were then fused by an unsupervised ensemble learning model, enabling the evaluation of the remaining combinations of AI model and SP method against the HIs evaluation criteria to determine the optimal final combination.
Figure 3: Framework for HI generation and evaluation — Step 1: Data Pre-Processing, Step 2: Signal Processing, Step 3: Feature Representation, Step 4: Feature Fusion — with prognostic criteria ($Mo$, PrPr, TrTr).
@@ 3.3 HIs criteria
Given the lack of true HI labels, evaluating the quality of the developed HIs necessitates leveraging known physics-based principles and facts. The HI criteria include $Mo$, PrPr, and $Tr$, and they are standard for this application [moradi2024designing, Moradi2024, Moradi2023Intelligent, Frusque2023, Baptista2022]. While these metrics do not directly measure physical damage, they evaluate whether the learned HI behaves in a physically meaningful and degradation-consistent manner.
MoMo measures the extent to which the HI shows only an increasing or decreasing trend. If maintenance is not performed, health should always decrease with an increasing number of cycles, and thus a monotonic HI should more closely model reality. The modified Mann-Kendall (MMK) formulation of $Mo$ is presented in Equation 1 [Moradi2024]:
@@
where $M$ is the number of specimens, $N$ is the number of GW measurements (timesteps), $y^{m}_{i}$ denotes an HI at measurement $i$ of specimen $m$, $t_{i}$ is similarly the time at the $i^{th}$ measurement, while $j$ can also represent a measurement number in this way. $sgn(\cdot)$ is the sign function. This formulation of $Mo$ is preferable as it considers time gaps greater than one unit, thus reducing the effect of noise and favoring monotonicity through the whole trend, as opposed to between adjacent points [Moradi2024].
PrPr measures the HIs’ deviation at failure between specimens. Preferably, all specimens should have the same value of HI at the end of life, such that they can be reliably used to predict failure. Equation 2 expresses $Pr$ [Moradi2024], where $std(\cdot)$ refers to the standard deviation between specimens.
@@
Finally, $Tr$ is a measure of the correlation between HIs of different units [Moradi2024]. The set of specimens should show the same HI pattern from beginning to end of life, as this makes their behaviour more predictable. $Tr$ is given by Equation 3, where $corr(\cdot)$ is the Pearson correlation function.
@@
In summary, $Mo$ quantifies how consistently the HI evolves monotonically under the assumption that health should not improve without maintenance, $Pr$ measures how closely specimens converge to a common HI value at failure, and $Tr$ evaluates whether specimens follow similar degradation trends throughout their lifetimes. To effectively evaluate a model’s performance on test specimens while mitigating bias towards highly matched training specimens and avoiding the obscuring of unmatched test specimen behaviour, alternative evaluation criteria must be established. These criteria should prioritise the behaviour of test specimens over those used during training, ensuring a more reliable assessment [Moradi2024, moradi2025novel]. These criteria include $Mo_{test}$, which evaluates the monotonicity of only the test specimen(s), and $Pr_{test}$, which assesses the prognosability by emphasizing the deviation of the test specimen(s) relative to those used for training. These metrics are defined in Equation 4 and Equation 5 [Moradi2024]:
@@
where the index traintrain or testtest represents train or test specimens, respectively, and thus $M^{train}$ refers to the number of the training specimens.
The fitness score $F_{all}$ is defined by Equation 6, and measures the overall quality of HIs [Moradi2024], while $F_{test}$ is defined similarly in Equation 7 with a focus on test units. In order to weight the criteria equally, the control constants $k_{Mo}$, $k_{Pr}$, and $k_{Tr}$ are set to equal 1. Both fitness scores $F_{all}$ and $F_{test}$ are calculated and discussed in this study.
@@
The prognostic criteria $Mo$, PrPr, and $Tr$ each take values in the range $[0,1]$, where 0 represents the worst behaviour and 1 represents the ideal expected behaviour of a degradation trajectory. The composite fitness scores $F_{all}$ and $F_{test}$ therefore lie in the range $[0,3]$, with 3 indicating the best overall performance across the three prognostic criteria. When reporting performance as a percentage (e.g. 92.3%), this corresponds to the fitness score normalised by its maximum value of 3.
@@ 3.4 Signal processing
SP techniques were applied to transform the measured signals into the frequency and time-frequency domains, in order to highlight relevant information, such as frequency components or noise isolation. These transformations enabled the extraction of relevant features for further analysis. The techniques used include FFT, STFT, EMD, and HT.
The FFT algorithm transforms the original temporal signal into the frequency domain by processing the entire time domain at once and outputting corresponding amplitudes for the underlying frequencies. However, it does not retain temporal variations in the signal, such as how these frequencies evolve over time, which may reveal additional information about the structure’s health. In contrast, STFT preserves time-frequency information by dividing the data into smaller time intervals and analyzing localised segments, making it advantageous for non-stationary signals where frequency content changes over time. The result of STFT is a 2D matrix, where time and frequency are the axes, and amplitude is the value at each position. Nevertheless, there is a nonlinear trade-off between time and frequency resolution: too many time segments reduces the available frequency samples, and vice versa. In this study, it was found that performing STFT with an FFT length of 250 maximised the number of data points, resulting in an overlap length of 125.
In the case of EMD, the sifting algorithm was implemented to iteratively decompose the signal into intrinsic mode functions (IMFs) in the time domain. The process involves constructing upper and lower envelopes of the signal using cubic splines to interpolate local maxima and minima, respectively. The mean of these envelopes is then subtracted from the signal at each iteration, refining the signal until it satisfies the conditions for an IMF. Suitable stopping conditions are necessary to avoid valuable information being lost [Boudraa2007]. Through trial and error, these criteria were set to a minimum standard deviation of $0.1$ and a maximum of 10 iterations.
The HT algorithm produces a complex-valued signal, the analytic signal, from a real signal. The real component represents the original signal, while the imaginary component is created by phase-shifting each frequency component by $\frac{\pi}{2}$ [Feldman2011]. The analytic signal therefore contains information concerning instantaneous amplitude and phase.
@@ 3.5 Feature extraction
Statistical features, as detailed in A, were extracted from the time, frequency, and time-frequency domains to characterise the signals comprehensively. The selected feature set aligns with established approaches in the literature [Moradi2023Intelligent]. They were chosen due to their proven capability to capture diverse and informative aspects of signal behaviour.
@@
For the first stage, features across different actuator-sensor paths and frequencies are averaged, resulting in 139 features per timestep (GW measurement) for each specimen. This averaging is intended to capture a robust representation of the specimen’s response, to emphasize consistent trends across multiple paths and frequencies whilst reducing the influence of local variations or noise. After obtaining all timesteps and specimens, the HI prognostic criteria defined in Section 3.3 are applied to evaluate each feature. In order to avoid anomalously poor features disproportionately impacting the SP method selected, the mean fitness score across all features is calculated to serve as a benchmark. Features with fitness scores exceeding this benchmark are retained for further analysis. This benchmark based filtering process acts as a feature selection step, comparable to Baraldi et al. [Baraldi2018OptimizationHI] method of identifying the most informative set of features to ensure that only consistently relevant features contribute to HI development. The SP method yielding features with the highest remaining mean fitness scores is identified as optimal for developing an efficient framework to extract HI, providing a systematic approach for feature selection rather then relying on arbitrary choices.
@@ 3.6 Feature fusion
Given the extracted and selected features as inputs, two deep learning models were developed to generate HIs, i.e. semi-supervised Diversity-DeepSAD and unsupervised DTC-VAE, which will be presented in this section.
@@ 3.6.1 Diversity-DeepSAD
The original DeepSAD model incorporates labeled training data in the loss function of the otherwise unsupervised Deep SVDD model [Ruff2019]. This approach transforms input data into a hyperspace of reduced dimensions, in which the distance to a point, the hypersphere centre $\mathbf{c}$, represents the extent to which the datapoint is an anomaly. Frusque et al. [Frusque2023] modified the loss function further to include a term to increase diversity, resulting in Diversity-DeepSAD for HI generation. This was demonstrated on a rotational machinery dataset with hundreds of measurement timesteps, a condition not often met in many other applications. For example, the ReMAP dataset may contain as few as 30 timesteps before failure. Consequently, it is necessary to explore the impact of alternative functions for generating auxiliary artificial labels and to evaluate the influence of the number of labeled measurements.
In prior studies [Ruff2019, Frusque2023], a binary set of auxiliary labels $\tilde{y}_{j}\in\{1,-1\}$ was employed, where a label of 1 was assigned to measurements representing a very healthy condition, and -1 corresponded to those reflecting a severely damaged state. This approach focuses solely on two extreme conditions. However, these two phases typically occupy only a small portion of the system’s entire lifetime, limiting the potential to fully leverage auxiliary labels. This issue is particularly pronounced in the severely damaged phase, where measurements are rare, often with only a single observation available at the point of failure for assigning the label -1. To address these limitations, this study generated continuous auxiliary labels $\tilde{y}_{j}\in[1,-1]$. Although many functions were considered, best performance was achieved through a simpler labeling scheme according to Equation 8:
@@
where a linear function was applied for the first and last quarters of the lifetime. The use of a linear function avoids imposing any assumed physical degradation which could reduce the flexibility of the model, given this is an extension from a pure binary $\pm 1$ formulation. The quarter ratio was established experimentally to produce improved results. As visualised in Figure 4a, the remaining data points do not leverage auxiliary labels, providing flexibility to the model, given the uncertainty of the true health status for these intermediate conditions.
It should be noted that auxiliary labels $\tilde{y}_{j}$ differ from the expected HI, denoted by $y_{j}$. Equation 8 therefore need not bear resemblance to the physical behaviour of the sample. Assuming the hypersphere centre $\mathbf{c}$ represents the healthy state, the HI is defined as the distance from that centre:
@@
where $z_{j}$ denotes the latent space representation obtained from the autoencoder. When $z_{j}$ is close to the healthy state $\mathbf{c}$, $y_{j}$ approaches 0, indicating minimal deviation from the healthy condition. Conversely, as $z_{j}$ moves farther from $\mathbf{c}$, $y_{j}$ increases, reflecting a more degraded condition. In this framework, the maximum value of 1 is assigned to $y_{j}$ in the fully damaged state. Consequently, the HI can freely exhibit linear or non-linear behaviour—with different main trends, such as exponential or polynomial behaviour—between these bounds, as illustrated in Figure 4b.
@@
Denoting the encoder function as $\phi$, the HI can be expressed as:
@@
where $\mathcal{L}_{L2}$ represents $L_{2}$ regularisation [gauss1809], minimising the sum of squares of node weights to encourage a greater distribution of smaller weights and prevent overfitting. $\nu$ and $\lambda$ are hyperparameters controlling the weights of the regularisation terms $\mathcal{L}_{L2}$ and $\mathcal{L}_{D}$, respectively. $\mathcal{L}_{D}$ is defined as follows:
$\mathcal{L}_{D}=\sum_{h=1}^{H}(\zeta_{h}-\ln(\zeta_{h}))$ (13)
where $\zeta_{h}$ represents the $h^{\text{th}}$ out of $H$ eigenvalues of the Gram matrix $G=Z^{T}Z$, in which $Z$ is the DeepSAD embedding. $\mathcal{L}_{D}$ promotes a greater distribution of training data in the model output [Frusque2023]. These three terms therefore work together to ensure that the model effectively learns the latent representations, while the regularisation terms help prevent overfitting and promote diversity within the learned embeddings.
A standard AE is always used to pretrain the network weights (learnable parameters $\theta$). This AE is formed by adding a decoder $\psi$ which inverts the Diversity-DeepSAD network layers, considering the mean squared difference between input and output data as a loss function, $\mathcal{L}_{AE}$. This loss function is used to pre-train the autoencoder, but is disregarded in the training of Diversity-DeepSAD in Equation 12. The hypersphere centre $\mathbf{c}$ is initialised by computing the average output from the model during a forward pass on the training data. Initially, $\mathbf{c}$ is set to a zero matrix. In the pretraining phase, denoted by $\phi_{\theta}^{[0]}$, the model performs a forward pass on the training data, and the outputs for each sample are accumulated. The number of samples is incremented with each data point, and the cumulative outputs are added to $\mathbf{c}$. After processing all the data, the matrix $\mathbf{c}$ is normalised by the total number of samples to obtain the average. To avoid numerical issues, any component of $\mathbf{c}$ with an absolute value smaller than $\epsilon$ is adjusted as follows:
@@
This adjustment ensures stability in the hypersphere centre values, helping the model to establish a reference point for the healthy state. The hypersphere centre $\mathbf{c}$ serves as a critical foundation for anomaly detection and HI estimation. The model architecture is illustrated in Figure 5, where red, yellow, and green represent pretraining, training, and model output, respectively.
@@
Diversity-DeepSAD architecture and hyperparameter optimisation: The Diversity-DeepSAD model was implemented with 6 hidden layers, each containing half the number of neurons as the previous layer, and leaky rectified linear units (Leaky ReLU) used as activation functions between layers. The hyperspace was set to be 16-dimensional (resulting in $\mathbf{c}$ with dimensions $4\times 4$), which was found to enhance the performance of AE during pretraining.
Hyperparameters with the strongest influence on model performance, i.e. batch size, learning rate (AE and DeepSAD stages) and the number of training epochs, were optimised via Bayesian optimisation. Optimisation was performed independently for every pair of frequency and feature type. Each run comprised 20 objective evaluations: the first 10 corresponded to uniform random sampling of the search space, and the remaining 10 were selected by a Gaussian-process surrogate model. The surrogate used a Matérn kernel with automatic relevance determination of the length scales, and a default acquisition strategy, which adaptively balances expected improvement, probability of improvement and lower confidence bound. The scalar optimisation objective was the $F_{all}$ score.
At every Bayesian optimisation iteration, a DeepSAD model was trained on the four training specimens of the fold, and $F_{all}$ was computed using only these units. The held-out specimen was excluded entirely during optimisation and used solely for testing. Across frequencies, optimisation trajectories typically stabilised after 15-18 evaluations, with flat incumbent curves and repeated sampling of similar hyperparameters, indicating convergence. For each (frequency, method), the hyperparameter set achieving the maximum observed $F_{all}$ was selected and retrained once on the full training set prior to evaluation on the test specimen.
All other hyperparameters were fixed based on literature [Frusque2023]: $\nu=10$, $\eta=10$, $\lambda=0.001$ and $\epsilon=1\cdot 10^{-6}$. Section 3.7, Section 4.6 and B discuss the sensitivity analysis to demonstrate that these fixed weights lie within broad regions of high performance, confirming their robustness.